OpenStudy (anonymous):

Use properties and definitions of operations to show that the statement is true. Justify each step. 15 x (3 / b) = 45 / b

OpenStudy (lgbasallote):

hmm \[15 \times (\frac 3b) \implies \frac{15 \times 3}{b}\]does that help?

OpenStudy (anonymous):

I'm supposed to do like a proof, I'll show you what I have

OpenStudy (anonymous):

(15 x 3) div (15 x b) = 45 div b : Distributive prop. of mult. and add. 45 div (15b) = 45 div b : Def. of mult.

OpenStudy (lgbasallote):

this isnt distributive property..

OpenStudy (anonymous):

I'm sorry I am just so confused....

OpenStudy (lgbasallote):

i think it's just multiplication....

OpenStudy (anonymous):

But the properties...?

OpenStudy (lgbasallote):

although there may be a proper term...im not familiar with fancy terms

OpenStudy (anonymous):

Do you want my list of properties? I could do this if I didnt have to write it as a proof.

OpenStudy (lgbasallote):

sure give me a list of the proeprties

OpenStudy (lgbasallote):

oh wait...i think i got it

OpenStudy (lgbasallote):

15 x(3/b) = (15x3)/b what i did was associative property

OpenStudy (anonymous):

Oh okay!

OpenStudy (anonymous):

Okay I think I got it thanks!

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

Do you think you can help me with another question similar to this?

OpenStudy (lgbasallote):

sure why not

OpenStudy (anonymous):

Okay thank you so much! (a+b) - c = a + (b-c)

OpenStudy (lgbasallote):

that's associative property too

OpenStudy (anonymous):

It would just be associative property right?

OpenStudy (anonymous):

Oh.. haha Thanks(:

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Thank you sooo much!

OpenStudy (lgbasallote):

welcome ^_^